# NJL-Chiral Soliton and the Nucleon Equation of State at supra-saturation density: Impact of Chiral Symmetry Restoration

### v1

@@ V Acknowledgment
The authors acknowledge the support of the CNRS-IN2P3 MAC masterproject, the project RELANSE ANR-23-CE31-0027-01 of the French National Research Agency (ANR), the European Union’s Horizon 2020 research and innovation program under grant agreement STRONG–2020-No824093. This work makes use of NumPy [98], SciPy [99], Matplotlib [100], jupyter [101] software packages, the Gnu Scientific Library software and reference manual [102].
@@ Appendix A Error Analysis
In our numerical implementation of the relaxation method, the possible sources of error arise from two main factors: (i) discretizations due to finite differences, and (ii) by the root-finding algorithm that minimizes the residuals (i.e., the difference between the left-hand side and right-hand side) of the field equations in eqs. 23, 24 and 25. The errors on the numerical solutions $y=[F(r),G(r),\omega(r)]$ due to the minimization procedure can be estimated using the upper bound error formula from Ref. [103]:
@@
where $\epsilon\sim\text{Ftol}\sim\text{tol}$ is the stopping criterion tolerance, $||\cdot||$ denotes the 2-norm, $J$ is the Jacobian matrix, and $e_{i}$ is the $i^{\text{th}}$ column of the identity matrix. The evolution of $\sigma_{i}$ for $N=500$, $\epsilon=\text{tol}=10^{-9}$, and boundary at $r_{\infty}=4$ fm using the parameter set of Ref. [58] is shown in fig. 17.
The discretizations error arises from evaluating the residuals of the equations using central finite differences, which introduce a local truncation error of $\sim\mathcal{O}(h^{2})$ in the derivatives $F^{\prime}(r),G^{\prime}(r)$, and $\omega^{\prime}(r)$. For linear problems, the global error remains $\sim\mathcal{O}(h^{2})$ [104, 105]. However, for nonlinear coupled systems, this error behavior is not always guaranteed.
To estimate the discretization error and to compare with a higher-accuracy solution, we employ Richardson extrapolation [105]. The error in the numerical solution $y_{h}(r)$ with step size $h$ is approximated by
@@
The absolute errors in the field profiles and their relative errors with respect to the improved solutions, computed within the extrapolation approach for the parameters of Ref. [58], are shown in figs. 18(a) and 18(b), respectively. For comparison, the $h^{2}$ behavior is also displayed in fig. 18(a). Using $N=500$, $r_{\infty}=4$ fm, and $h\sim 10^{-2}$ fm, the numerical errors in the model observables are found to be:
@@
Figure 19: Dependence of soliton properties on $r_{\infty}$ for the relaxation method. Upper Panel: Variation of $\sqrt{\langle r_{B}^{2}\rangle}$. Middle panel: $\sqrt{\langle r_{S}^{2}\rangle}$. Lower Panel: Variation of soliton mass $M_{\rm slt}$ for different choices of $r_{\infty}$. All results correspond to the soliton properties obtained with meson parameters of Ref. [58].
@@ Appendix B Nucleon-Scalar coupling $g_{s}$ and vector properties of the NJL model
Following the definition of eq. 18, the soliton mass $M_{s}(s)$ can be evolved by allowing both scalar-sector parameters ($F_{\pi},M_{\pi}$) and vector-sector parameters ($g_{v},M_{v}$) to vary with $s$. In this self-consistent treatment, the nuclear scalar coupling $g_{s}$ receives positive contributions from the scalar sector and negative contributions from the vector sector, leading to a reduced overall value.
In contrast, several works, see Refs. [77, 78], incorporate in-medium chiral symmetry restoration only through the scalar properties ($M_{\pi},F_{\pi},M_{\sigma}$), effectively treating the scalar field as the sole mediator of restoration while keeping the vector sector fixed at its vacuum values. For clarity, we distinguish two cases:
@@
    Case 1: mass of the soliton $M_{\rm slt}(s)$ evaluated following  section II, with $M_{v}$ and $g_{v}$ fixed at their vacuum values (solid line in fig. 20).
@@
    Case 2: mass of the soliton $M_{\rm slt}(s)$ evaluated following  section II with both $M_{v}(s)$ and $g_{v}(s)$ evolving with $s$ (dashed line in fig. 20).
@@ Appendix C In-Medium Meson Properties
We display in-medium meson properties in nuclear matter for different NJL parameterizations in fig. 21. For a given baryon density $\rho_{N}$, we first determine the in-medium nuclear scalar field $s$ (equivalently ${\cal S}(s)$) using eq. 51, as described in section III.2.2. From this, the corresponding in-medium meson properties $M_{\pi}({\cal S})$, $F_{\pi}({\cal S})$, $M_{\sigma}({\cal S})$, and $M_{v}({\cal S})$ are obtained via eqs. 7, 8, 9 and 10, following the method outlined in section II. For clarity, all quantities are normalized to their vacuum values, so that the plots highlight their relative changes with density. As the density increases and chiral symmetry is progressively restored ($|s|\to f_{\pi}$), the in-medium pion mass $M_{\pi}({\cal S})$ approaches the in-medium scalar mass $M_{\sigma}({\cal S})$.
@@
Figure 21: Relative changes of in-medium meson properties, normalized to their vacuum values, as functions of nuclear matter density ($\rho_{N}$) and the corresponding nuclear scalar field $s$.

### v2

@@ 5 Acknowledgment
The authors acknowledge the support of the CNRS-IN2P3 MAC masterproject, the project RELANSE ANR-23-CE31-0027-01 of the French National Research Agency (ANR), the European Union’s Horizon 2020 research and innovation program under grant agreement STRONG–2020-No824093. This work makes use of NumPy vanderWalt:2011bqk , SciPy Virtanen:2019joe , Matplotlib Hunter:2007 , jupyter jupyter software packages, the Gnu Scientific Library software and reference manual galassi2018scientific .
@@ Appendix A Error Analysis
In our numerical implementation of the relaxation method, the possible sources of error arise from two main factors: (i) discretizations due to finite differences, and (ii) by the root-finding algorithm that minimizes the residuals (i.e., the difference between the left-hand side and right-hand side) of the field equations in Eqs. (23)-(25). The errors on the numerical solutions $y=[F(r),G(r),\omega(r)]$ due to the minimization procedure can be estimated using the upper bound error formula from Ref. Keller1974 :
@@
where $\epsilon\sim\text{Ftol}\sim\text{tol}$ is the stopping criterion tolerance, $||\cdot||$ denotes the 2-norm, $J$ is the Jacobian matrix, and $e_{i}$ is the $i^{\text{th}}$ column of the identity matrix. The evolution of $\sigma_{i}$ for $N=500$, $\epsilon=\text{tol}=10^{-9}$, and boundary at $r_{\infty}=4$ fm using the parameter set of Ref. Fukushima2020 is shown in Fig. 18.
The discretizations error arises from evaluating the residuals of the equations using central finite differences, which introduce a local truncation error of $\sim\mathcal{O}(h^{2})$ in the derivatives $F^{\prime}(r),G^{\prime}(r)$, and $\omega^{\prime}(r)$. For linear problems, the global error remains $\sim\mathcal{O}(h^{2})$ kelley1995iterative ; Richard2010 . However, for nonlinear coupled systems, this error behavior is not always guaranteed.
To estimate the discretization error and to compare with a higher-accuracy solution, we employ Richardson extrapolation Richard2010 . The error in the numerical solution $y_{h}(r)$ with step size $h$ is approximated by
@@
The absolute errors in the field profiles and their relative errors with respect to the improved solutions, computed within the extrapolation approach for the parameters of Ref. Fukushima2020 , are shown in Fig. 19(a) and 19(b), respectively. For comparison, the $h^{2}$ behavior is also displayed in Fig. 19(a). Using $N=500$, $r_{\infty}=4$ fm, and $h\sim 10^{-2}$ fm, the numerical errors in the model observables are found to be:
@@
Figure 20: Dependence of soliton properties on $r_{\infty}$ for the relaxation method. Upper Panel: Variation of $\sqrt{\langle r_{B}^{2}\rangle}$. Middle panel: $\sqrt{\langle r_{S}^{2}\rangle}$. Lower Panel: Variation of soliton mass $M_{\rm slt}$ for different choices of $r_{\infty}$. All results correspond to the soliton properties obtained with meson parameters of Ref. Fukushima2020 .
@@ Appendix B Nucleon-Scalar coupling $g_{s}$ and vector properties of the NJL model
Following the definition of Eq. (18), the soliton mass $M_{s}(s)$ can be evolved by allowing both scalar-sector parameters ($F_{\pi},M_{\pi}$) and vector-sector parameters ($g_{v},M_{v}$) to vary with $s$. In this self-consistent treatment, the nuclear scalar coupling $g_{s}$ receives positive contributions from the scalar sector and negative contributions from the vector sector, leading to a reduced overall value.
In contrast, several works, see Refs. CHRISTOV1990 ; Alkofer1991 , incorporate in-medium chiral symmetry restoration only through the scalar properties ($M_{\pi},F_{\pi},M_{\sigma}$), effectively treating the scalar field as the sole mediator of restoration while keeping the vector sector fixed at its vacuum values. For clarity, we distinguish two cases:
@@
    Case 1: mass of the soliton $M_{\rm slt}(s)$ evaluated following Section 2, with $M_{v}$ and $g_{v}$ fixed at their vacuum values (solid line in Fig. 21).
@@
    Case 2: mass of the soliton $M_{\rm slt}(s)$ evaluated following Section 2 with both $M_{v}(s)$ and $g_{v}(s)$ evolving with $s$ (dashed line in Fig. 21).
@@ Appendix C In-Medium Meson Properties
We display in-medium meson properties in nuclear matter for different NJL parameterizations in Fig. 22. For a given baryon density $\rho_{N}$, we first determine the in-medium nuclear scalar field $s$ (equivalently ${\cal S}(s)$) using Eq. (51), as described inSection 3.2.2. From this, the corresponding in-medium meson properties $M_{\pi}({\cal S})$, $F_{\pi}({\cal S})$, $M_{\sigma}({\cal S})$, and $M_{v}({\cal S})$ are obtained via Eqs. (7), (8),  (9), and (10), respectively, following the method outlined in Section 2. For clarity, all quantities are normalized to their vacuum values, so that the plots highlight their relative changes with density. As the density increases and chiral symmetry is progressively restored ($|s|\to f_{\pi}$), the in-medium pion mass $M_{\pi}({\cal S})$ approaches the in-medium scalar mass $M_{\sigma}({\cal S})$.
@@
Figure 22: Relative changes of in-medium meson properties, normalized to their vacuum values, as functions of nuclear matter density ($\rho_{N}$) and the corresponding nuclear scalar field $s$. We have also shown evolution of the ratio among the pion and the scalar meson masses with density.
Appendix D Effect of Boundary Conditions for the Soliton in Dense Matter
As discussed in Section 3.2.2, the density evolution of the soliton is driven by the scalar mean field $\cal{S}$ generated by the surrounding medium, which is treated as uniform in the mean-field approximation. In symmetric nuclear matter, the pion and $\rho$-meson mean fields vanish outside the soliton, so their boundary conditions remain identical to the vacuum case. The $\omega$ meson, however, asymptotically can acquire a nonzero in-medium expectation value (see Ref. Li2024 ).
In uniform matter with baryon density $\rho_{N}$, the $\omega$ mean field satisfies
$\bar{\omega}=\frac{g_{v}}{M_{v}^{2}}\,\rho_{N},$ (63)
and the soliton profile may asymptotically approach
$\omega_{\rm soliton}(r)\big|_{r\to r_{\infty}}=-\bar{\omega}.$ (64)
This prescription corresponds to treating the soliton as a localized excitation embedded in infinite dense matter, with boundary conditions imposed at $r_{\infty}\to\infty$, as discussed for instance in Ref. Li2024 . We refer to this choice as B.C-2, while the vacuum-like boundary conditions used throughout the main text are denoted as B.C-1. Explicitly,
B.C-1: $\omega^{\prime}(0)=0,\quad\omega(r_{\infty}\to\infty)=0,$ (65)
B.C-2: $\omega^{\prime}(0)=0,\quad\omega(r_{\infty}\to\infty)=-\bar{\omega}(\rho_{N}),$ (66)
with identical conditions for the $F$ and $G$ profiles in both cases.
We have verified that implementing B.C-2 does not lead to any significant modification of the soliton core properties. In particular, the soliton-based equation of state  Eq. (57), the EoS extracted from the soliton center, the density evolution of the hard-core radius, and the truncation density of soliton based EoS remain essentially unchanged, as shown in Fig. 23.
Figure 23: (a) EoSs obtained under different B.Cs are displayed. For reference, the evolution of the central pressure ($p_{\rm slt}^{c}$) as a function of the central energy density ($\varepsilon_{\rm slt}^{c}$) of the soliton is also shown. (b) Evolution of the hard-core radius as a function of density obtained under different B.Cs. The comparison is performed for the NJL-F parameterization.
For completeness and to comment, we also explored an alternative construction inspired by Wigner-Seitz treatments of dense matter Nagai2006 , where the soliton is placed at the center of a spherical unit cell of radius
$R_{\rm WS}=\left(\frac{4\pi}{3}\rho_{N}\right)^{-1/3},$ (67)
and boundary conditions are imposed at $r_{\infty}=R_{\rm WS}$. We denote this choice as B.C-3, with
$\omega^{\prime}(0)=0,\quad\omega(R_{\rm WS})=-\bar{\omega}(\rho_{N}).$ (68)
Representative meson profiles and energy-pressure distributions obtained under B.C-3 are shown in Fig. 24 and Fig. 25 respectively. In this case, we find that the stability condition $\int d^{3}r\,p(r)=0$ is violated once $R_{\rm WS}\lesssim 1~\mathrm{fm}$, indicating that this prescription is not compatible with a stable soliton solution in the present NJL-based framework. A thorough investigation of such modified boundary conditions would require a separate study, similar to analyses performed for Skyrmion crystals or solitons on a lattice.
Figure 24: Meson profiles of the soliton at different bulk baryon densities under B.C-3. The comparison is performed for the NJL-F parameterization.
Figure 25: Distribution of the soliton’s energy density and pressure at different bulk nuclear matter densities. The comparison is performed for the NJL-F parameterization.
We therefore conclude that the boundary conditions adopted in the main text (B.C-1) are the most consistent choice within the mean-field and Born-Oppenheimer approximations employed in this work. Furthermore, a similar treatment for the scalar field profile could be implemented in future work by going beyond the Born-Oppenheimer approximation. In such an extended framework, the soliton would receive additional medium-induced contributions from both scalar and vector mesons, together with consistently modified boundary conditions, allowing for a more complete description of in-medium soliton dynamics.